# From Equations to Insights: Unraveling Symbolic Structures in PDEs with LLMs

### v2

@@ 1 Introduction
Numerous grand challenges in scientific fields, including physics, engineering, biology, and chemistry, can be modeled as partial differential equations (PDEs), which relate unknown functions defined on a multidimensional domain (e.g., in terms of space and time) to its rates of change with respect to each of those variables [11]. Mathematically, a PDE takes the following form
$Du=f,\quad\text{in }\Omega,\quad Bu=g,\quad\text{on }\partial\Omega,$ (1)
where $D$ denotes a differential operator defining the dynamics, $Bu=g$ represents the boundary condition (e.g., Dirichlet, Neumann, and Robin), and $\Omega\subseteq\mathbb{R}^{d}$ is the domain with boundary $\partial\Omega$.
@@
Before solving a PDE, one must understand certain fundamental properties of the solution $u$ that follow from the PDE itself and from the data, namely, $f$ and $g$. These considerations align with the classical theory of PDEs [11], which focuses on establishing well-posedness (i.e., the existence and uniqueness of solutions) and their regularity and continuity. However, the underlying symbolic relationship among $u$, $f$ and $g$ often remains underexplored. To address this gap, we pose the following question:
Assuming that the PDE (1) admits an analytical solution $u$, and given the operators in $f$ and $g$, how are the operators in $u$ related to those in $f$ and $g$?
@@
$L(u)=\|Du-f\|^{2}_{L^{2}(\Omega)}+\lambda\|Bu-g\|^{2}_{L^{2}(\partial\Omega)},$ (2)
@@
$u^{*}=\mathrm{argmin}\left\{L(u)\;:\;u\in L^{2}(\Omega)\right\}.$ (3)
@@
To uncover the symbolic relationship, we leverage large language models (LLMs) as predictive tools. The process begins with the generation of a comprehensive, structured dataset composed of symbolic expressions derived from a diverse range of PDE types, including elliptic, parabolic, and hyperbolic equations, along with their associated boundary and initial conditions. These symbolic expressions are represented as computational trees (see Figure 5), where nodes are populated with unary (e.g., sin, cos, exp) and binary operators (e.g., +, -, *, /). This tree structure captures the hierarchical nature of mathematical expressions and ensures a systematic representation of the PDE solutions.
@@
To prepare the data for LLM processing, the symbolic expressions are tokenized and converted into postfix notation (Reverse Polish Notation). This format is chosen because it aligns well with the sequential input requirements of LLMs and eliminates the need for parentheses, simplifying the parsing process. More importantly, it reflects the relative positional relationship of operators in the expression. The tokenized dataset is then used to fine-tune foundational LLMs such as T5, BART, and Llama. During fine-tuning, the models are trained to predict the sequence of operators that constitute the symbolic solution to a new PDE. The training objective is to maximize the accuracy of operator sequence prediction, ensuring that the model can reliably identify the minimal set of operators required to represent the true solution.
@@
Once fine-tuned, the LLM acts as a predictive tool to generate operator sets for new PDE problems. The predicted operator sets are highly refined, containing only the operators necessary to construct the solution, thereby eliminating redundant or irrelevant operators. This refinement is crucial for the subsequent application of the finite-expression method (FEX), a reinforcement learning based symbolic regression technique that searches for the optimal combination of operators and constants to approximate the PDE solution. By providing FEX with a reduced and targeted operator set, the combinatorial search space is significantly narrowed. This reduction not only enhances the accuracy of the solution by focusing on relevant operators, but also dramatically improves computational efficiency by reducing the number of potential combinations that need to be evaluated.
@@
In summary, the integration of LLMs into this workflow transforms the process of symbolic PDE solving. By generating structured datasets, fine-tuning models to predict operator sequences, and leveraging these predictions to guide FEX, we achieve a more accurate and computationally efficient approach to discovering symbolic solutions for complex PDEs. This methodology bridges the gap between machine learning and symbolic mathematics, offering great potential for scientific computing and engineering applications. Particularly, our contributions are summarized as follows:
@@
(1) We formally pose the problem of discovering the symbolic relationship between the solution of a PDE and the associated problem data, including the boundary conditions and right-hand-side functions. Establishing this connection can provide deeper insights into the underlying structure of PDEs, enabling more precise analytical solutions and facilitating the development of efficient numerical methods which provides high-quality analytical solutions. By highlighting the potential impact of such relationships, our work lays the foundation for further studies in this area, aiming to bridge the gap between problem formulation and solution representation.
(2) We propose fine-tuning LLMs to learn the symbolic relationship between problem data and the solution of a PDE, demonstrating remarkable effectiveness in predicting the operators that characterize the solution. By integrating this approach with the FEX method, our numerical findings show that while LLMs may not be inherently suited for directly solving PDEs or other complex scientific problems, they provide substantial benefit in identifying underlying structures and symbolic patterns. This capability significantly streamlines and enhances the solution process. Our work underscores the potential of LLMs as transformative tools in the analytical and computational study of PDEs, bridging the gap between symbolic reasoning and sophisticated mathematical problem-solving.
@@ 2 Theoretical Insight for Poisson Equations
$-\Delta u=f\quad\text{in }\Omega,\qquad u=g\quad\text{on }\partial\Omega,$ (4)
@@
    $\tilde{u}$ is constructed using only the elementary arithmetic operators +,×$+,\;\times,\;(\cdot)^{2}$ and the operators appearing in the evaluation of $\mathcal{D}(\cdot)$, $f(\cdot)$, and $g(\cdot)$,
@@
Proof.
See Appendix B. ∎
This theorem offers a partial constructive resolution to the open research question posed in the previous section: it identifies the specific operators needed to represent a $\delta$-accurate approximation of the PDE solution $u$ in terms of the operators used to define the source term $f$, the boundary condition $g$, and the domain geometry via the distance function $\mathcal{D}$. The central insight is that an accurate approximation $\tilde{u}$ can be constructed using only a limited set of algebraic operations, along with the computational building blocks inherent in $f$, $g$, and $\mathcal{D}$. This result carries several important implications. (1) In contexts where one seeks symbolic models for physical phenomena, the theorem guarantees the existence of a low-complexity symbolic approximation of $u$ that relies solely on a structured and interpretable set of primitive operations. (2) It also informs the design of neural architectures for PDE approximation, suggesting that imposing operator priors, i.e., by reusing the operators present in $f$ and $g$, is sufficient to achieve arbitrarily high approximation accuracy, thereby promoting efficiency and interpretability. (3) The polynomial bound on computational complexity with respect to $\delta^{-1}$ underscores the practical feasibility and scalability of such approximations. Taken together, these insights lay a foundation for a broader theory of operator-preserving approximations for general PDEs, with significant potential impact in scientific machine learning, computational physics, and algorithmic model design.
3 Methodology
3.1 Binary Computational Trees for Expressions
Expressions in PDEs are represented as binary computational trees, as depicted in Figure 5 in Appendix A. These trees provide a structured and hierarchical way to encode mathematical expressions, enabling efficient exploration and manipulation of the solution space.
Each node in the computational tree encodes an operator, which may be either a unary or binary operator. Unary Operators: Unary operators act on a single operand and include a variety of nonlinear transformations such as trigonometric functions ($\sin(\cdot)$, $\cos(\cdot)$, $\tan(\cdot)$), exponential functions ($\exp(\cdot)$), logarithmic functions ($\lg(\cdot)$, $\ln(\cdot)$), and other elementary operations (e.g., $\sqrt{\cdot}$, ||$\cdot$|| ⋅ |, and power functions $(\cdot)^{k}$ for $k\in\mathbb{Z}$). Each unary operator node is augmented with two scalar parameters, $\alpha$ and $\beta$, which enable the composition of the operator with an affine transformation, yielding expressions of the form $\alpha\cdot u(\cdot)+\beta$, where $u(\cdot)$ denotes the unary operation. This parameterization allows the tree to flexibly represent both nonlinear and linearly adjusted transformations, thereby enhancing its expressiveness. Binary Operators: Binary operators operate on two operands and include the standard arithmetic operations: addition ($+$), subtraction (--), multiplication ($\times$), and division (///). Each binary operator node combines the outputs of its left and right child subtrees to produce more complex expressions. For example, a multiplication node may combine two distinct sub-expressions, thereby enabling the construction of richer symbolic forms through hierarchical composition.
3.2 Data Generation for Equation Types
Figure 1: Data generation pipeline.
3.3 Fine-tuning LLM
With a generated data set, we can then fine-tune a certain LLM to learn the symbolic relastion in PDEs. The whole pipeline is summarized in Figure 2, and we will describe each component in detail for the rest of this subsection.
Figure 2: Overview of the fine-tuning pipeline
$\underbrace{\texttt{Type:}<\mathrm{PDE}>\,\texttt{|}\,\texttt{RHS:}<\mathrm{ops}>\,\texttt{|}\,\texttt{BC\_Type:}<\mathrm{bc}>\,\texttt{|}\,\texttt{Solution:}}_{\text{PROMPT}}\;\|\|\underbrace{\texttt{<op}_{1}\ \dots\ \texttt{op}_{n}\,\texttt{<EOS>}\texttt{>}}_{\text{TARGET}}$
3.4 Inferencing
Since the model’s raw output is a token sequence, a post-processing pipeline is required to extract a clean and interpretable set of predicted operators. The first step involves identifying the unique set of operators by removing duplicate tokens from the generated sequence. Following this, we perform error checking to detect and eliminate misspelled or malformed tokens that may result from tokenization artifacts or character-level prediction errors (issues that can arise because the model learns statistical patterns rather than enforcing strict syntactic correctness). Such errors are particularly likely when dealing with rare or domain-specific mathematical symbols. Any invalid or unrecognized tokens are subsequently discarded to ensure the integrity of the final operator set.
To quantitatively evaluate the model’s predictive accuracy, we compute the squared $\ell_{2}$ norm of the difference between the predicted and ground-truth operator sets. Specifically, we encode each operator set as a binary vector over a fixed dictionary of $n$ possible operators, where each vector component indicates the presence or absence of a given operator. This encoding transforms the comparison into a well-defined vector space problem. An illustrative example of this binary encoding scheme can be found in Appendix C.1. Then, the squared distance between two such operator sets, represented by two binary vectors $y\in\mathbb{R}^{n}$ and $z\in\mathbb{R}^{n}$, can be defined as $\|y-z\|^{2}=\sum_{i=1}^{n}\left(y_{i}-z_{i}\right)^{2}.$ Clearly, it measures the number of mismatched operators between the two operator sets.
4 Experimental Results
4.1 Experimental Setup
[…]
4.2 Effectiveness of Fine-tuned LLMs
In this subsection, we evaluate the performance of the fine-tuned BART, T5 and Llama3 models in predicting operator sets for the two PDE types. Specifically, we track the average number of mismatched operators on the test dataset for each epoch. The computational results are summarized in Figure 3. Interested readers are referred to the Table 3 in Appendix C.2 for the detailed computational results.
4.3 Effectiveness of LLM-informed FEX
In this section, we compare the performance of LLM-informed FEX with the fine-tuned LLaMA-8B model and the (original) uninformed FEX. To this end, we evaluated the performance of the FEX algorithm on a set of 100 randomly generated PDE examples, including 50 Poisson equations and 50 Linear Conservation Law equations. Since FEX is a randomized reinforcement learning approach, each instance was solved five times to obtain average computational results, including the number of iterations until convergence, computation time, and solution error. The detailed computational results can be found in Appendix C.3, and Table 1 and Figure4 summarize the statistics of these results.
PDE Type Metric Uninformed FEX LLM-Informed FEX Speedup
Conservation Avg. Iter 28.52 5.23 5.45$\times$
Avg. Time (m) 23.67 3.95 6.00$\times$
Poisson Avg. Iter 28.73 6.47 4.44$\times$
Avg. Time (m) 62.07 14.09 4.41$\times$
Table 1: Comparison on computational efficiency.
The computational results show that the LLM-informed FEX significantly improves efficiency over the uninformed FEX. A key advantage is the reduction in the number of operators used. The uninformed FEX always relies on a fixed set of 3 binary and 9 unary operators. In contrast, the LLM-informed FEX dynamically and effectively selects smaller and more relevant operator sets. This smalle and accurate search space leads to a substantial drop in the number of iterations needed for convergence. This efficiency gain further results in significant computational time savings. Despite this excellent speedup, the accuracy remains competitive: when the uninformed FEX converges, both methods produce similar approximation errors. And the LLM-informed FEX demonstrates better performance in terms of stablility as shown in Figure 4. These findings demonstrate that integrating the fine-tuned LLM into symbolic machine learning based PDE solvers enables more efficiency and reliability via informed operator selection.
5 Conclusions
Acknowledgments and Disclosure of Funding
The authors were partially supported by the US National Science Foundation under awards DMS-2244988, DMS-2206333, the Office of Naval Research Award N00014-23-1-2007, and the DARPA D24AP00325-00.
Appendix A Binary Expression Trees
The building blocks of the FEX methods in [30] are illustrated in the following Table 5. By increasing the depth of the rexpression tree, the number of functions that can be represented can be significantly increased. Also, the mechanism can automatically handle high-dimensional variables.
Figure 5: Computational expression tree [30].
Appendix B Proof of Theorem 1
@@
[…]
@@
[…]
@@
$f^{(M,\{K_{i}\},J)}(x):=$ $\;\frac{1}{MJ}\sum_{i=1}^{M}\sum_{k=1}^{M_{i}}\sum_{j=1}^{J}\mathcal{D}\left(\bar{X}_{k}(x,Y_{i,1},\dots,Y_{i,k})\right)^{2}$
×$\qquad\times f\left(\bar{X}_{k-1}(x,Y_{i,1},\dots,Y_{i,k-1})+\mathcal{D}(\bar{X}_{k-1}(x,Y_{i,1},\dots,Y_{i,k-1}))y_{i,j,k}\right),$
@@
[…]
@@
if one chooses suitable $M,\;J$ and $K_{i},\;i=1,\dots,M$ all polynomial in $\delta^{-1}$ [15]. Therefore, the proof is completed.
Appendix C Detailed Computational Results
C.1 Illustrative Example for Binary Encoding
Consider the following two functions:
$h_{1}(x_{1},x_{2}):=(5x_{1})^{2}+\sin(3x_{2})\cdot x_{2},\quad h_{2}(x_{1}):=5\exp(2x_{1})\cdot(\cos(6x_{1}))^{3}.$
The binary representation of them is provided in Table 2.

### v3

@@ Abstract
keywords:
Large Language Models, Finite Expression Method, High Dimensional PDEs, Operator Relation
{MSCcodes}
65N75, 68T20, 90C15
@@ 1 Introduction
Partial differential equations (PDEs) are the mathematical language of the natural sciences: they encode how fields evolve and interact by linking an unknown function on space–time to its derivatives. From turbulent flows and elastic structures to reaction–diffusion in biology and quantum systems, many grand scientific challenges are governed by PDE models [12].
Mathematically, a strong-form PDE constrains an unknown function $u:\Omega\to\mathbb{R}$ on a domain $\Omega\subseteq\mathbb{R}^{d}$ by relating $u$ to its partial derivatives together with boundary data
(1) $Du=f,\quad\text{in }\Omega,\quad Bu=g,\quad\text{on }\partial\Omega.$
Here $D$ is a differential operator of order $k$ that may depend on $x$, $u$, and derivatives of $u$ up to order $k$; a linear prototype is $Du=\sum_{|\alpha|\leq k}a_{\alpha}(x)\,D^{\alpha}u$ with multi-index $\alpha=(\alpha_{1},\ldots,\alpha_{d})$ and $D^{\alpha}u:=\partial^{|\alpha|}u/\partial x_{1}^{\alpha_{1}}\cdots\partial x_{d}^{\alpha_{d}}$, while $f:\Omega\to\mathbb{R}$ is a given source term and $g:\partial\Omega\to\mathbb{R}$ is given. The boundary operator $B$ prescribes data on $\partial\Omega$. Some commonly used boundary conditions include $Bu=\gamma_{0}u=u|_{\partial\Omega}=g$ (Dirichlet), $Bu=\gamma_{1}u=\partial_{\boldsymbol{n}}u=g$ with $\boldsymbol{n}$ the outward unit normal (Neumann), or $Bu=\gamma_{1}u+\kappa\,\gamma_{0}u=g$ for $\kappa\geq 0$ (Robin) where $\gamma_{0}:H^{1}(\Omega)\to H^{1/2}(\partial\Omega)$ and $\gamma_{1}:H^{1}(\Omega)\to H^{-1/2}(\partial\Omega)$ denote continuous trace maps. For time-dependent problems on $Q=(0,T)\times\Omega$, one augments the system with the initial condition $u(0,\cdot)=u_{0}$ in $\Omega$. Classical examples include the parabolic and hyperbolic models
$u_{t}-\nabla\!\cdot(A\nabla u)=f,\qquad u_{tt}-\nabla\!\cdot(A\nabla u)=f,$
where $A$ denotes the diffusion/conductivity tensor. When classical derivatives are unavailable, the statement $Du=f$ and $Bu=g$ is interpreted in a weak (variational) sense on Sobolev spaces, where existence and uniqueness follow under standard boundedness and coercivity assumptions (e.g., by the Lax–Milgram theorem). And the principal symbol of $D$ then classifies the PDE as elliptic, parabolic, or hyperbolic.
@@
Before solving a PDE, one must understand certain fundamental properties of the solution $u$ that follow from the PDE itself and from the data, namely, $f$ and $g$. These considerations align with the classical theory of PDEs [12], which focuses on establishing well-posedness (i.e., the existence and uniqueness of solutions) and their regularity and continuity. However, the underlying symbolic relationship among $u$, $f$ and $g$ often remains underexplored. To address this gap, we pose the following question: Assuming that the PDE (1) admits an analytical solution $u$, and given the operators in $f$ and $g$, how are the operators in $u$ related to those in $f$ and $g$?
@@
(2) $L(u)=\|Du-f\|^{2}_{L^{2}(\Omega)}+\lambda\|Bu-g\|^{2}_{L^{2}(\partial\Omega)},$
@@
(3) $u^{*}=\mathrm{argmin}\left\{L(u)\;:\;u\in L^{2}(\Omega)\right\}.$
@@
To uncover the symbolic relationship, we leverage large language models (LLMs) as predictive tools. The process begins with the generation of a comprehensive, structured dataset composed of symbolic expressions derived from a diverse range of PDE types, including elliptic, parabolic, and hyperbolic equations, along with their associated boundary and initial conditions. These symbolic expressions are represented as computational trees (see Fig. 1), where nodes are populated with unary (e.g., sin, cos, exp) and binary operators (e.g., +, -, *, /). This tree structure captures the hierarchical nature of mathematical expressions and ensures a systematic representation of the PDE solutions.
Figure 1: Computational expression tree [32].
@@
To prepare the data for LLM processing, the symbolic expressions are tokenized and converted into postfix notation (Reverse Polish Notation). This format is chosen because it aligns well with the sequential input requirements of LLMs and eliminates the need for parentheses, simplifying the parsing process. More importantly, it reflects the relative positional relationship of operators in the expression. The tokenized dataset is then used to fine-tune foundational LLMs. During fine-tuning, the models are trained to predict the sequence of operators that constitute the symbolic solution to a new PDE. The training objective is to maximize the accuracy of operator sequence prediction, ensuring that the model can reliably identify the minimal set of operators required to represent the true solution.
@@
Once fine-tuned, the LLM acts as a predictive tool to generate operator sets for new PDE problems. The predicted operator sets are highly refined, containing only the operators necessary to construct the solution, thereby eliminating redundant or irrelevant operators. This refinement is crucial for the subsequent application of the finite-expression method (FEX), a reinforcement learning–based symbolic regression technique that searches for the optimal combination of operators and constants to approximate the PDE solution. By providing FEX with a reduced and targeted operator set, the combinatorial search space is significantly narrowed. This reduction not only enhances the accuracy of the solution by focusing on relevant operators, but also dramatically improves computational efficiency by reducing the number of potential combinations that need to be evaluated. While we adopt FEX in this work, it is not the only viable approach. If the symbolic relationship between a PDE and its solution can be effectively learned, then a wide range of symbolic regression methods [26], such as evolutionary algorithms [38], can be directly applied using the learned operator set.
@@
In summary, the integration of LLMs into this workflow transforms the process of symbolic PDE solving. By generating structured datasets, fine-tuning models to predict operator sequences, and leveraging these predictions to guide PDE solvers, we achieve a more accurate and computationally efficient approach to discovering symbolic solutions for complex PDEs. This methodology bridges the gap between machine learning and symbolic mathematics, offering great potential for scientific and engineering applications. Particularly, our contributions are summarized as follows:
  1. 1.
@@
    PDE symbolic relations. We formally pose the problem of discovering the symbolic relationship between the solution of a PDE and the associated problem data, including the boundary conditions and source functions. Establishing this connection can provide deeper insights into the underlying structure of PDEs, enabling more precise analytical solutions and facilitating the development of efficient numerical methods which provides high-quality analytical solutions. By highlighting the potential impact of such relationships, our work lays the foundation for further studies, aiming to bridge the gap between problem formulation and solution representation.
  2. 2.
    LLM-guided symbolic PDE solvers. We fine-tune LLMs to predict the operator set that captures the solution’s structure and couple these predictions with the FEX pipeline to search over symbolic programs that produce fully interpretable surrogates $u_{\theta}$. Experiments show that, while LLMs are not direct PDE solvers, their operator- and geometry-aware priors sharply prune the search over expressions, improving FEX’s sample efficiency and stability and yielding higher-quality closed-form or near–closed-form solutions across PDE families. The same operator-selection front end is modular and can be integrated with other symbolic PDE solvers where operator choice is the critical bottleneck.
  3. 3.
    Theory for policy-gradient–based symbolic learning. We provide convergence guarantees for a stochastic projected policy-gradient method that maximizes the physics-informed return over symbolic programs. In particular, under standard conditions, we prove that projected updates converge to first-order stationary policies and establish a sample complexity of $\mathcal{O}(\epsilon^{-4})$ to reach an $\epsilon$-stationary point in expectation, thereby placing the LLM+FEX pipeline on firm theoretical footing among state-of-the-art alternatives.
@@ 2 Theoretical Insight for Poisson Equations
(4) $\Delta u=f\quad\text{in }\Omega,\qquad u=g\quad\text{on }\partial\Omega,$
@@
    $\tilde{u}$ is constructed using only the elementary arithmetic operators $\times,\;(\cdot)^{2}$ and the operators appearing in the evaluation of $\mathcal{D}(\cdot)$, $f(\cdot)$, and $g(\cdot)$,
@@
Proof 2.2.
@@
(5) $\;u(x)=\mathbb{E}_{x}\left[g(X_{\tau(\Omega)})\right]+\frac{1}{2}\mathbb{E}_{x}\left[\sum_{k\geq 1}\mathcal{D}(\bar{X}_{k})^{2}E_{0}\right],$
where
$E_{0}:=\mathbb{E}_{0}\left[\int_{0}^{\tau(B(0,1))}f\left(\bar{X}_{k-1}+D(\bar{X}_{k-1})\cdot t\right)dt\right].$
@@
$\sqrt{\int_{\Omega}\left(\mathbb{E}_{x}\left[g(X_{\tau(\Omega)})\right]-g^{(M,\{K_{i}\})}(x)\right)^{2}dx}\leq\delta.$
@@
$f^{(M,\{K_{i}\},J)}(x)=\frac{1}{MJ}\sum_{i=1}^{M}\sum_{k=1}^{M_{i}}\sum_{j=1}^{J}\mathcal{D}\left(\bar{X}_{k}(x,Y_{i,1},\dots,Y_{i,k})\right)^{2}\times f\left(Z_{k-1}\right),$
where
$Z_{k-1}=\bar{X}_{k-1}(x,Y_{i,1},\dots,Y_{i,k-1})+\mathcal{D}(\bar{X}_{k-1}(x,Y_{i,1},\dots,Y_{i,k-1}))y_{i,j,k},$
@@
$\sqrt{\int_{\Omega}\left(\mathbb{E}_{x}\left[\sum_{k\geq 1}\mathcal{D}(\bar{X}_{k})^{2}E_{0}\right]-f^{(M,\{K_{i}\},J)}(x)\right)^{2}dx}\leq\delta,$
where
$E_{0}:=\mathbb{E}_{0}\left[\int_{0}^{\tau(B(0,1))}f\left(\bar{X}_{k-1}+\mathcal{D}(\bar{X}_{k-1})\cdot t\right)dt\right],$
@@
if one chooses suitable M,JM,\;J and $K_{i},\;i=1,\dots,M$ all polynomial in $\delta^{-1}$ [16]. Therefore, the proof is completed.
The objective of the above theorem is to establish a proof-of-concept framework to guide symbolic discovery of PDE solutions using modern machine learning approaches, including LLMs. This theorem offers a partial constructive resolution to the open research question posed in the previous section: it identifies the specific operators needed to represent a $\delta$-accurate approximation of the PDE solution $u$ in terms of the operators used to define the source term $f$, the boundary condition $g$, and the domain geometry via the distance function $\mathcal{D}$. The central insight is that an accurate approximation $\tilde{u}$ can be constructed using only a limited set of algebraic operations, along with the computational building blocks inherent in $f$, $g$, and $\mathcal{D}$. This result carries several important implications. (1) In contexts where one seeks symbolic models for physical phenomena, the theorem guarantees the existence of a low-complexity symbolic approximation of $u$ that relies solely on a structured and interpretable set of primitive operations. (2) It also informs the design of neural architectures for PDE approximation, suggesting that imposing operator priors, i.e., by reusing the operators present in $f$ and $g$, is sufficient to achieve arbitrarily high approximation accuracy, thereby promoting efficiency and interpretability. (3) The polynomial bound on computational complexity with respect to $\delta^{-1}$ underscores the practical feasibility and scalability of such approximations. Taken together, these insights lay a foundation for a broader theory of operator-preserving approximations for general PDEs, with significant potential impact in scientific machine learning, computational physics, and algorithmic model design. We plan to pursue a unified analysis for more general classes of PDEs in future work, and our current numerical results on linear conservation laws already provide clear evidence supporting this direction.
3 Methodology
3.1 Binary Computational Trees for Expressions
Expressions in PDEs are represented as binary computational trees, as depicted in Figure 1. These trees provide a structured and hierarchical way to encode mathematical expressions, enabling efficient exploration and manipulation of the solution space.
Each node in the computational tree encodes an operator, which may be either a unary or binary operator. Unary Operators: Unary operators act on a single operand and include a variety of nonlinear transformations such as trigonometric functions ($\sin(\cdot)$, $\cos(\cdot)$, $\tan(\cdot)$), exponential functions ($\exp(\cdot)$), logarithmic functions ($\lg(\cdot)$, $\ln(\cdot)$), and other elementary operations (e.g., $\sqrt{\cdot}$, $|\cdot|$, and power functions $(\cdot)^{k}$ for $k\in\mathbb{Z}$). Each unary operator node is augmented with two scalar parameters, $\alpha$ and $\beta$, which enable the composition of the operator with an affine transformation, yielding expressions of the form $\alpha\cdot u(\cdot)+\beta$, where $u(\cdot)$ denotes the unary operation. This parameterization allows the tree to flexibly represent both nonlinear and linearly adjusted transformations, thereby enhancing its expressiveness. Binary Operators: Binary operators operate on two operands and include the standard arithmetic operations: addition (++), subtraction (-), multiplication ($\times$), and division (//). Each binary operator node combines the outputs of its left and right child subtrees to produce more complex expressions. For example, a multiplication node may combine two distinct sub-expressions, thereby enabling the construction of richer symbolic forms through hierarchical composition.
3.2 Data Generation for Equation Types
Figure 2: Data generation pipeline.
3.3 Fine-tuning LLM
With a generated data set, we can then fine-tune a certain LLM to learn the symbolic relation in PDEs. The whole pipeline is summarized in Figure 3, and we will describe each component in detail for the rest of this subsection.
Figure 3: Overview of the fine-tuning pipeline
$\;\underbrace{\texttt{Type:}<\mathrm{PDE}>\,\texttt{|}\,\texttt{RHS:}<\mathrm{ops}>\,\texttt{|}\,\texttt{BC\_Type:}<\mathrm{bc}>\,\texttt{|}\,\texttt{Solution:}}_{\text{PROMPT}}$
$\;\underbrace{\texttt{<op}_{1}\ \dots\ \texttt{op}_{n}\,\texttt{<EOS>}\texttt{>}}_{\text{TARGET}}$
In this work, we intentionally adopted a simple and minimal prompt design to isolate the core capabilities of the fine-tuned language model and avoid introducing prompt-specific biases. However, incorporating explicit task descriptions or more structured prompts could potentially enhance performance, and exploring such prompt engineering strategies is a valuable research direction.
3.4 Inferencing
Since the model’s raw output is a token sequence, a post-processing pipeline then extracts a clean and interpretable set of predicted operators. The first step involves identifying the unique set of operators by removing duplicate tokens from the sequence. Following this, we perform error checking to detect and eliminate misspelled or malformed tokens that may result from tokenization artifacts or character-level prediction errors (issues that can arise because the model learns statistical patterns rather than enforcing strict syntactic correctness). Such errors are particularly likely when dealing with rare or domain-specific mathematical symbols. Any invalid or unrecognized tokens are subsequently discarded to ensure the integrity of the final operator set.
To measure the model’s predictive accuracy, we compute the squared $\ell_{2}$-norm of the difference between the predicted and ground-truth operator sets. Specifically, we encode each operator set as a binary vector over a fixed dictionary of $n$ possible operators, where each vector component indicates the presence or absence of a given operator. This encoding transforms the comparison into a well-defined vector space problem. An illustrative example of this binary encoding scheme with
$h_{1}(x_{1},x_{2}):=(5x_{1})^{2}+\sin(3x_{2})\cdot x_{2},\quad h_{2}(x_{1}):=5\exp(2x_{1})\cdot(\cos(6x_{1}))^{3},$
is showed in Table 1. Then, the squared distance between two such operator sets, represented by two binary vectors $y\in\mathbb{R}^{n}$ and $z\in\mathbb{R}^{n}$, can be defined as $\|y-z\|^{2}=\sum_{i=1}^{n}\left(y_{i}-z_{i}\right)^{2}.$ Clearly, it measures the number of mismatched operators between the two operator sets. The resulted mismatch can then be computed as $7$, indicating that the two functions are very different from each other.